the loaded roller coaster has a mass of 2500 kg. the starting point of the run is 22m above point A; the run from B to C is 120m long; the run from C to D is 100m. at D, brakes are applied, bringing the coaster to a halt 15m further along a E. if the friction from B to D is 0, find

(a) the potential energy of the roller coaster at the starting point

(b) KE of the roller coaster at C

(c) the speed of the roller coaster at D; the average braking force applied between D and E


1
Expert's answer
2022-01-16T13:09:05-0500

a)

"E_A=mgh_A=0.55~MJ,"

b)

"K_C=E_A-E_C=mg(h_A-h_C)=0.18~MJ,"

c)

"v_D=\sqrt{\frac{2E_D}m}=\sqrt{\frac{2mg(h_A-h_D)}m}=\sqrt{2g(h_A-h_D)}=11.8~\frac ms,"

d)

"Fl_{DE}=E_D=mg(h_A-h_D),"

"F=\frac{mg(h_A-h_D)}{l_{DE}}=23.3~kN."